Together a chair, a table, and a lamp cost $562. The chair costs 4 times as much as the lamp, and the table costs $23 less than the chair. Calculate the cost of the table, the chair, and the lamp.

Answer by rfer(16322)

chair=4x
lamp=x
table=4x-23
4x+x+4x-23=562
9x=585
x=585/9
x=65
4x=260
4x-23=237